# Solution of some tiling open problems of Propp, Lai, and some related results

### v2

@@ Abstract.
As another application, we obtain a semi-factorization result for the number of lozenge tilings of a large class of hexagonal regions with holes (obtained by starting with an arbitrary symmetric hexagon with holes, and translating all the holes one unit lattice segment in the same direction). This in turn leads to the solution of two open problems posed by Lai, to an extension of a result due to Fulmek and Krattenthaler, and to exact enumeration formulas for some new families of hexagonal regions with holes.
@@ 1. Introduction
Two particular classes of results of this kind concern honeycomb graphs with holes along a symmetry axis (for some illustrative examples, see [5]), and Aztec rectangles with holes along the axis of symmetry (see e.g. [4]). As it turns out, in both these cases interesting results hold also when the collection of holes is translated some distance away from the symmetry axis. Indeed, Krattenthaler proved such formulas for Aztec rectangles with unit holes (see [21]). Also, Lai conjectured (see [24]) that simple product formulas exist for the number of lozenge tilings of two families of regions, both obtained from symmetric hexagons with certain types of holes along the symmetry axis, by translating all the holes one unit lattice segment in the same direction [^1] By identifying lozenge tilings of a region on the triangular lattice with perfect matchings of its planar dual, this is equivalent to enumerating perfect matchings of a honeycomb graph with holes.: one family corresponds to the holes being an arbitrary collection of like-oriented triangles of side-length two along the symmetry axis (the untranslated version of this is one of the families of regions for which a simple product formula is proved in [5]), while in the other there is a single hole in the shape of a shamrock, symmetric about the symmetry axis (the original, untranslated case is treated in [7] and [25], where a simple product formula is given for it). However, in both these situations the factorization theorem cannot be applied directly, because the dual graph of the involved regions are not quite symmetric.
@@
In Section 3 we use Theorem 2.1 to solve four open problems of Propp on the enumeration of trimer tilings on the hexagonal lattice.
@@ 2. Two general results
Figure 1. Left. A plane bipartite symmetric graph $G$. Right. The graphs $G^{+}$ (on top) and $G^{-}$ (on the bottom) corresponding to $G$.
@@
The weight of a matching $\mu$ is defined to be the product of the weights of the edges contained in $\mu$. The matching generating function of a weighted graph $G$, also denoted333 When all weights are 1, this becomes just the number of perfect matchings of $G$. by $\operatorname{M}(G)$, is the sum of the weights of all perfect matchings444 For weighted graphs, $\operatorname{M}_{v}(G)$ denotes the sum of the weights of all $v$-near-perfect matchings of $G$. of $G$. The matching generating function is clearly multiplicative with respect to disjoint unions of graphs. We will henceforth assume that all graphs under consideration are connected.
Let $G$ be a weighted symmetric graph that is also bipartite. For definiteness, choose the leftmost vertex on the symmetry axis $\ell$ to be white. We define two subgraphs $G^{+}$ and $G^{-}$ as follows.
Given a vertex $u$ of $G$ on the symmetry axis, we call the operation of deleting all edges above $\ell$ which are incident to $u$ cutting above $u$; similarly, we call deleting all edges below $\ell$ incident to $u$ cutting below $u$. Perform cutting operations above all white $a_{i}$’s and black $b_{i}$’s and below all black $a_{i}$’s and white $b_{i}$’s. Note that this procedure yields cuts of the same kind at the endpoints of each edge lying on $\ell$ (see the proof of Theorem 2.1(a) for a justification). Reduce the weight of each such edge by half; leave all other weights unchanged.
As shown in the proof of Theorem 2.1(a), the graph produced by the above procedure is disconnected into one part lying above and one lying below $\ell$. Denote the portion above $\ell$ by $G^{+}$, and the one below $\ell$ by $G^{-}$; see the picture on the right in Figure 1 for an illustration of this procedure (the edges whose weight has been reduced by half are marked by 1/2).
Part (a) of the following result is a slight strengthening of the original factorization theorem of [4, Theorem 1.2] (namely, we show that the assumption made there that the graph is separated by its symmetry axis can be dropped). The new version is contained in part (b).
@@
((((a)))). If $G$ has an even number of vertices, then
@@
((((b)))). Suppose $G$ has an odd number of vertices and $v$ is a vertex on the unbounded face of $G$ lying off the symmetry axis. Assume (without loss of generality, as $G$ is symmetric) that $v$ is above the symmetry axis. Then
@@
Figure 2. The graph $\overline{G}$ corresponding to the graph $G$ shown in Figure 1 with one vertex removed (since the graph in Figure 1 has the same number of white and black vertices, we have to delete from it a vertex on $\ell$ of the color opposite to the color of $v$ in order to obtain a graph meeting the hypotheses of part (b) of Theorem 2.1).
@@
To prove (a), it suffices to show that our assumptions imply that the graph $G_{0}$ produced by the cutting procedure described before the statement of the theorem is disconnected into one part lying above and one lying below $\ell$ (i.e., that the graph is separated by its symmetry axis, in the terminology of [4, Theorem 1.2]); indeed, all the other assumptions in the latter are met by hypothesis. Namely, we need to show that there exists a Jordan arc $J$ in the plane connecting some point $A$ on $\ell$ to some point $B$ on $\ell$ — where $A$ is to the left of $a_{1}$, and $B$ is to the right of $b_{\operatorname{w}(G)}$ — so that the portion $H$ of $G_{0}$ which is above or on $\ell$ is above $J$, while the portion $K$ of $G_{0}$ which is below or on $\ell$ is below $J$.
@@
Secondly, note that $G_{0}$ cannot have any edge $\{u,v\}$ with $u$ above $\ell$ and $v$ below $\ell$. Indeed, if $u$ and $v$ are not mirror images with respect to $\ell$, the symmetry of $G$ implies that the mirror images $u^{\prime}$ and $v^{\prime}$ of $u$ and $v$ form another edge $\{u^{\prime},v^{\prime}\}$ of $G_{0}$; but then the edges $\{u,v\}$ and $\{u^{\prime},v^{\prime}\}$ would cross, which is a contradiction, since $G$ is a plane graph. On the other hand, if $u$ and $v$ are mirror images of each other, let $P$ be a path in $G$ connecting $u$ to $a_{1}$ (as $G$ is assumed to be connected, such a path exists). Then the mirror image $P^{\prime}$ of $P$ across $\ell$ is a path in $G$ connecting $v$ to $a_{1}$. Since $P$ and $P^{\prime}$ have the same length and $G$ is bipartite, it follows that $u$ and $v$ have the same color, so the edge $\{u,v\}$ has endpoints of the same color; this contradicts the fact that $G$ is bipartite.
By the second to last paragraph, we can connect $A$ to $B$ by a Jordan arc $J$ which does not cross any edge of $G_{0}$ incident to a vertex on $\ell$. By the previous paragraph, this arc $J$ does not cross any other edge of $G_{0}$ either. This proves our claim, and completes the proof of part (a).
@@
To prove part (b), note first that unless the difference between the number of white and black vertices of $G$ is ±$\pm 1$, and in addition $v$ has the color of the majority of vertices, equation (2.2) holds trivially, as both sides equal zero. Therefore, without loss of generality we may assume that these conditions hold.
@@
Next, define $\overline{G}$ to be the graph obtained from $G$ by including a new vertex $w$ on $\ell$ contained in the unbounded face of $G$, and the two new edges $\{w,v\}$ and $\{w,v^{\prime}\}$ (both weighted by 1), where $v^{\prime}$ is the mirror image of $v$ (see Figure 2). Then $\overline{G}$ is a plane, bipartite, weighted symmetric graph, to which part (a) of the theorem can be applied.
@@
If $v$ is white, $w$ is black, and (since $w$ is also a $b_{i}$-type vertex) the cutting operation at $w$ prescribed by the factorization theorem occurs above $w$. Therefore, $\overline{G}^{+}$ is precisely $G^{+}$. Furthermore, in $\overline{G}^{-}$ vertex $w$ has degree one, and after removing the edge $\{u,v^{\prime}\}$ that is forced to match it in every perfect matching, the resulting graph (which has the same matching generating function as $\overline{G}^{+}$, as the removed edge has weight 1) is precisely $G^{-}\setminus v^{\prime}$. Thus the factorization theorem applied to $\overline{G}$ yields
@@
Draw the triangular lattice so that one family of lattice lines is horizontal, and let $H$ be a non-degenerate (i.e., with all sides of positive length) hexagonal region on this lattice, symmetric about a vertical symmetry axis555 We will be interested in symmetric (or nearly-symmetric) regions obtained from $H$ by making in it some holes, and since the symmetry of such regions seems to be more readily appreciated visually when the symmetry axis is vertical, we adopt this point of view for the second result in this section. $\ell$. Let $R_{0}$ be the region obtained from $H$ by removing an arbitrary collection ${\mathcal{C}}$ of unit triangles, so that ${\mathcal{C}}$ is symmetric about $\ell$ and contains no unit triangles touching the southern or southeastern edge of $H$. Define $R$ to be the region obtained from $R_{0}$ by translating all unit triangles in ${\mathcal{C}}$ one unit in the southeastern lattice direction. We call the region $R$ a nearly symmetric hexagon with holes. Therefore, while the outer, hexagonal boundary of the region $R$ is symmetric with respect to the original symmetry axis $\ell$, the holes in $R$ are symmetric about the line $\ell^{\prime}$, the translation of $\ell$ half a unit to the right.
Let $G$ be the planar dual graph of $R$. Even though $G$ is not symmetric about $\ell^{\prime}$, let us consider the graphs $G^{+}$ and $G^{-}$ obtained from it by rotating it 90 degrees counterclockwise and performing cutting operations around the vertices on $\ell^{\prime}$ as prescribed by the factorization theorem (assuming, in particular, that after the rotation the leftmost vertex on $\ell^{\prime}$ is white). In addition to these two graphs, consider also the graphs $\widehat{G}^{+}$ and $\widehat{G}^{-}$ obtained from $G$ by performing cutting operations around the vertices on $\ell^{\prime}$ as prescribed by the factorization theorem, but assuming this time that after the rotation the leftmost vertex on $\ell^{\prime}$ is black. Define $R^{+}$, $R^{-}$, $\widehat{R}^{+}$ and $\widehat{R}^{-}$ to be the lattice regions whose planar duals are the graphs $G^{+}$, $G^{-}$, $\widehat{G}^{+}$ and $\widehat{G}^{-}$, respectively666 If $R$ is the region on the triangular lattice whose planar dual is the graph $G$, and $G$ has weights on its edges, the lozenge position in $R$ corresponding to any given edge of $G$ comes weighted with the weight of that edge..
@@
Let $R$ be a nearly symmetric hexagon with holes, and assume that the symmetric region $R_{0}$ from which $R$ was constructed777 I.e. the region obtained from $R$ by translating all the holes one unit lattice segment in the northwest direction. admits at least one lozenge tiling. Then the number of lozenge tilings888 We denote the matching generating function of the lozenge tilings of a region $R$ (in which lozenge positions may carry weights) by $\operatorname{M}(R)$, because these lozenge tilings can be identified with perfect matchings of the (weighted) planar dual graph of the region $R$. of $R$ is given by
@@
$\operatorname{M}(\hat{R}\setminus\{a,d\})\operatorname{M}(\hat{R}\setminus\{b,c\})=\operatorname{M}(\hat{R})\operatorname{M}(\hat{R}\setminus\{a,b,c,d\})+\operatorname{M}(\hat{R}\setminus\{a,c\})\operatorname{M}(\hat{R}\setminus\{b,d\}).$ (2.5)
Note that after removing some forced lozenges, $\hat{R}\setminus\{a,d\}$ is precisely the region $R$ whose tilings we want to enumerate (this is in fact what guided our construction of the extended region $\hat{R}$), and $\hat{R}\setminus\{b,c\}$ is a symmetric region (see the bottom left picture in Figure 4). Furthermore, each of the two regions $\hat{R}\setminus\{a,b,c,d\}$ and $\hat{R}\setminus\{a,c\}$ can be viewed as being obtained from a symmetric region with a unit dent on the boundary, by removing some forced lozenges (see the bottom center picture and top right picture, respectively, in Figure 4). The remaining two regions $\hat{R}$ and $\hat{R}\setminus\{b,d\}$ are also equivalent to symmetric regions with a unit defect on the boundary, but we first need to enlarge them to see this. In the case of $\hat{R}$, we add one layer of unit triangles on the top right side of $\hat{R}$, and remove the bottom-most of these (see the top center picture in Figure 4). The region $\hat{R}\setminus\{b,d\}$ can be handled similarly, but we first remove from it the forced lozenges along the top and bottom sides (see the bottom right picture in Figure 4).
Therefore, viewing the regions in (2.5) as described above, we can apply Theorem 2.1(a) to $\hat{R}\setminus\{b,c\}$, and Theorem 2.1(b) to the four regions $\hat{R}$, $\hat{R}\setminus\{a,b,c,d\}$, $\hat{R}\setminus\{a,c\}$, and $\hat{R}\setminus\{b,d\}$. When applying [^9] Strictly speaking, we apply Theorem 2.1 to their planar dual graphs, and then consider the regions whose planar duals are the resulting graphs. Theorem 2.1, the powers of two in the factorizations of $\operatorname{M}(\hat{R}\setminus\{b,c\})$, $\operatorname{M}(\hat{R})$, and $\operatorname{M}(\hat{R}\setminus\{a,c\})$ are $2^{w_{l^{\prime}}(R)}$, while those in the factorizations of $\operatorname{M}(\hat{R}\setminus\{a,b,c,d\})$ and $\operatorname{M}(\hat{R}\setminus\{b,d\})$ are $2^{w_{l^{\prime}}(R)-1}$. Figure 5 shows how the five regions are split via Theorem 2.1 (the shaded ellipses indicate lozenge positions weighted by $1/2$); note that the first branch of Theorem 2.1(b) is applied when the unit boundary dent (which corresponds to $v$) has the same orientation as the top unit triangle along the symmetry axis (which corresponds to $a_{1}$), and the second branch when the orientations are opposite. One readily sees that the resulting “half”-subregions — we call them parts — have the following properties:
@@
$(i)$ The left part of $\hat{R}\setminus\{b,c\}$ is the same (up to forced lozenges) as the left part of $\hat{R}\setminus\{a,b,c,d\}$, and the right part of $\hat{R}\setminus\{b,c\}$ is the same (up to some forced lozenges) as the right part of $\hat{R}$.
$(ii)$ The left part of $\hat{R}\setminus\{b,c\}$ is the same as the right part of $\hat{R}\setminus\{b,d\}$, and the right part of $\hat{R}\setminus\{b,c\}$ is the same (up to forced lozenges) as the left part of $\hat{R}\setminus\{a,c\}$.
$(iii)$ The remaining two parts of $\hat{R}$ and $\hat{R}\setminus\{a,b,c,d\}$ (the ones which are not the same as any of the parts of $\hat{R}\setminus\{b,c\}$) can be identified with the two parts of $R$ obtained when we split $R$ using one of the two zigzag lines along $\ell^{\prime}$, as prescribed by the factorization theorem (compare the two picture at the center in Figure 5 and the left picture in Figure 6).
$(iv)$ The remaining two parts of $\hat{R}\setminus\{a,c\}$ and $\hat{R}\setminus\{b,d\}$ (the ones which cannot be identified with any part of $\hat{R}\setminus\{b,c\}$), can be identified (up to forced lozenges) with the two parts of $R$ when we split $R$ using the other zigzag line along $\ell^{\prime}$, as prescribed by the factorization theorem (compare the two pictures on the right in Figure 5 and the right picture in Figure 6).
By applying Theorem 2.1 to each of the five regions in (2.5) besides $\hat{R}\setminus\{a,d\}$ (which, as we pointed out, after removing the forced lozenges becomes precisely the region $R$ whose tilings we want to enumerate, equation (2.5) becomes an equality in which the left hand side is the product of three factors, while the right hand side is the sum of two products, each having four factors.
By observations $(i)$ and $(ii)$ above, both factors on the left hand side besides $\operatorname{M}(\hat{R}\setminus\{a,d\})$ also appear in each of the two terms on the right hand side. By assumption, $\hat{R}\setminus\{b,c\}$ (which is the same as the region $R_{0}$) is tileable; hence, by Theorem 2.1, both these factors are non-zero. Dividing by them, the left hand side of the equation discussed in the previous paragraph becomes just $\operatorname{M}(R)$. Furthermore, by $(iii)$, the two remaining factors in the first term on the right hand side are precisely $\operatorname{M}(R^{+})$ and $\operatorname{M}(R^{-})$. Finally, by $(iv)$, the two remaining factors in the second term on the right hand side are precisely $\operatorname{M}(\widehat{R}^{+})$ and $\operatorname{M}(\widehat{R}^{-})$. Thus, the equation at which we have arrived is precisely (2.4), and the proof is complete.

### v3

# Propp’s benzels and Lai’s nearly symmetric hexagons with holes

@@ Abstract.
As another application, we obtain a semi-factorization result for the number of lozenge tilings of a large class of hexagonal regions with holes (obtained by starting with an arbitrary symmetric hexagon with holes, and translating all the holes one unit lattice segment in the same direction). This in turn leads to the solution of two open problems posed by Lai and to an extension of a result due to Fulmek and Krattenthaler, which results in exact enumeration formulas for some new families of hexagonal regions with holes.
@@ 1. Introduction
Two particular classes of results of this kind concern honeycomb graphs with holes along a symmetry axis (for some illustrative examples, see [5]), and Aztec rectangles with holes along the axis of symmetry (see e.g. [4]). As it turns out, in both these cases interesting results hold also when the collection of holes is translated some distance away from the symmetry axis. Indeed, Krattenthaler proved such formulas for Aztec rectangles with unit holes (see [21]). Also, Lai conjectured (see [24]) that simple product formulas exist for the number of lozenge tilings of two families of regions, both obtained from symmetric hexagons with certain types of holes along the symmetry axis, by translating all the holes one unit lattice segment in the same direction [^1] By identifying lozenge tilings of a region on the triangular lattice with perfect matchings of its planar dual, this is equivalent to enumerating perfect matchings of a honeycomb graph with holes.: one family corresponds to the holes being an arbitrary collection of like-oriented triangles of side-length two along the symmetry axis (the untranslated version of this is one of the families of regions for which a simple product formula is proved in [5]), while in the other there is a single hole in the shape of a shamrock, symmetric about the symmetry axis (the original, untranslated case is treated in [7] and [25], where a simple product formula is given for it). However, in both these situations the factorization theorem cannot be applied directly, because the dual graphs of the involved regions are not quite symmetric.
@@
In Section 3 we use Theorem 2.1 to solve four open problems of Propp (Problems 8 through 11 in [30]) on the enumeration of trimer tilings333 More precisely, these are tilings by triples of contiguous unit hexagons, which in the dual picture become trimer coverings. For ease or reference we call them simply trimer tilings. on the hexagonal lattice.
@@ 2. Two general results
All the graphs in this section will be assumed to be bipartite. The two classes of vertices will be called white and black.
@@
Figure 1. Illustrating Theorem 2.1(a): A plane bipartite symmetric graph $G$ (left), and the resulting graphs $G^{+}$ and $G^{-}$ (right).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2. Illustrating the case of Theorem 2.1(b) when the removed boundary vertex $v$ is white. A plane, bipartite symmetric graph $G$ with an odd number of vertices on the symmetry axis, with the number of white vertices being one more than the number of black vertices (top left); the corresponding graphs $G^{+}$ and $G^{-}$ (top right). When a white vertex $v$ is removed from the boundary (bottom left), the factorization involves the graphs $G^{+}$ and $G^{-}\setminus v^{\prime}$, where $v^{\prime}$ is the mirror image of $v$ (bottom right).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3. Illustrating the case of Theorem 2.1(b) when the removed boundary vertex $v$ is black. A plane, bipartite symmetric graph $G$ with an odd number of vertices on the symmetry axis, with one more black than white vertices (top left); the corresponding graphs $G^{+}$ and $G^{-}$ (top right). When a black vertex $v$ is removed from the boundary (bottom left), the factorization involves the graphs $G^{+}\setminus v$ and $G^{-}$ (bottom right).
@@
The weight of a matching $\mu$ is defined to be the product of the weights of the edges contained in $\mu$. The matching generating function of a weighted graph $G$, also denoted444 When all weights are 1, this becomes just the number of perfect matchings of $G$. by $\operatorname{M}(G)$, is the sum of the weights of all perfect matchings555 For weighted graphs, $\operatorname{M}_{v}(G)$ denotes the sum of the weights of all $v$-near-perfect matchings of $G$. of $G$. The matching generating function is clearly multiplicative with respect to disjoint unions of graphs. We will henceforth assume that all graphs under consideration are connected.
Let $G$ be a weighted symmetric graph that is also bipartite. For definiteness, choose the leftmost vertex on the symmetry axis $\ell$ to be white. We also assume that $G$ has at least one vertex on the symmetry axis $\ell$ (note that this implies that the symmetry is color-preserving; see the fifth paragraph of the proof of Theorem 2.1). We define two subgraphs $G^{+}$ and $G^{-}$ as follows.
Given a vertex $u$ of $G$ on the symmetry axis, we call the operation of deleting all edges above $\ell$ which are incident to $u$ cutting above $u$; similarly, we call deleting all edges below $\ell$ incident to $u$ cutting below $u$. Perform cutting operations above all white $a_{i}$’s and black $b_{i}$’s and below all black $a_{i}$’s and white $b_{i}$’s. Note that this procedure yields cuts of the same kind at the endpoints of each edge lying on $\ell$ (see the fourth paragraph in the proof of Theorem 2.1 for a justification). Reduce the weight of each such edge by half; leave all other weights unchanged.
As shown in the proof of Theorem 2.1(a), the graph $G_{0}$ produced by the above procedure is disconnected into one part lying above and one lying below $\ell$. Denote the portion above $\ell$ by $G^{+}$, and the one below $\ell$ by $G^{-}$; see the picture on the right in Figure 1 for an illustration of this procedure (the edges whose weight has been reduced by half are marked by 1/2).
Part (a) of the following result is a slight strengthening of the original factorization theorem of [4, Theorem 1.2] (namely, we show that the assumption made there that the graph is separated by its symmetry axis can be dropped). The new version is contained in part (b); its two cases are illustrated in Figures 2 and 3.
@@
((a)). If $G$ has an even number of vertices, then
@@
((b)). Suppose $G$ has an odd number of vertices and $v$ is a vertex on the unbounded face of $G$ lying off the symmetry axis. Assume ((without loss of generality, as $G$ is symmetric)) that $v$ is above the symmetry axis. Then
@@
(a) This part is almost exactly the same as the second author’s factorization theorem [4, Theorem 1.2]): the only difference is that one of the assumptions in the latter — namely, that the graph is separated by its symmetry axis (the meaning of this will be recalled below) — is not present in the statement of the current theorem. We show that this condition is in fact a consequence of the other assumptions (a fact overlooked at the time of writing of [4]). Then part (a) will follow directly from [4, Theorem 1.2].
The needed separation condition is that the graph $G_{0}$ produced by the cutting procedure described before the statement of the theorem is disconnected into one part lying above and one lying below $\ell$. We show below that this is a consequence of the current assumptions.
What we need to show is that there exists a Jordan arc $J$ in the plane connecting some point $A$ on $\ell$ to some point $B$ on $\ell$ — where $A$ is to the left of $a_{1}$, and $B$ is to the right of $b_{\operatorname{w}(G)}$ — so that the portion $H$ of $G_{0}$ which is above or on $\ell$ is above $J$, while the portion $K$ of $G_{0}$ which is below or on $\ell$ is below $J$.
@@
Secondly, note that $G_{0}$ cannot have any edge $\{u,v\}$ with $u$ above $\ell$ and $v$ below $\ell$. Indeed, if $u$ and $v$ are not mirror images with respect to $\ell$, the symmetry of $G$ implies that the mirror images $u^{\prime}$ and $v^{\prime}$ of $u$ and $v$ form another edge $\{u^{\prime},v^{\prime}\}$ of $G_{0}$; but then the edges $\{u,v\}$ and $\{u^{\prime},v^{\prime}\}$ would cross, which is a contradiction, since $G$ is a plane graph (see the picture on the left in Figure 4). On the other hand, if $u$ and $v$ are mirror images of each other, let $P$ be a path in $G$ connecting $u$ to $a_{1}$ (as $G$ is assumed to be connected, such a path exists). Then the mirror image $P^{\prime}$ of $P$ across $\ell$ is a path in $G$ connecting $v$ to $a_{1}$. Since $P$ and $P^{\prime}$ have the same length and $G$ is bipartite, it follows that $u$ and $v$ have the same color, so the edge $\{u,v\}$ has endpoints of the same color; this contradicts the fact that $G$ is bipartite.
The connected components of $G_{0}\cap\ell$ are either single vertices or paths with at least one edge. By the definition of $G_{0}$, all edges incident to a single-vertex component are confined to one side of $\ell$. Furthermore, by the second to last paragraph, all edges of $G_{0}$ incident to the vertices of a path-component are also contained in one of the two half-planes bounding $\ell$. Therefore we can connect $A$ to $B$ by a Jordan arc $J$ that stays close to $\ell$ and meanders between the connected components of $G_{0}\cap\ell$ so that it does not cross any edge of $G_{0}$ incident to a vertex on $\ell$. In fact, this arc $J$ does not cross any other edge of $G_{0}$ either. Indeed, as $J$ can stay arbitrarily close to $\ell$, such an edge of $G_{0}$ would need to have its endpoints on opposite sides of $\ell$, which we showed in the previous paragraph that is impossible. This proves our claim, and completes the proof of part (a).
Refer to caption Refer to caption
Figure 4. Left. Illustrating the argument in the fifth paragraph of the proof of Theorem 2.1. Right. The graph $\overline{G}$ corresponding to the graph $G$ shown in Figure 1 with one vertex removed (since the graph in Figure 1 has the same number of white and black vertices, we have to delete from it a vertex on $\ell$ of the color opposite to the color of $v$ in order to obtain a graph meeting the hypotheses of part (b) of Theorem 2.1).
@@
(b) To prove the second part, note first that unless the difference between the number of white and black vertices of $G$ is $\pm 1$, and in addition $v$ has the color of the majority of vertices, equation (2.2) holds trivially, as both sides equal zero. Therefore, without loss of generality we may assume that these conditions hold.
@@
Next, define $\overline{G}$ to be the graph obtained from $G$ by including a new vertex $w$ on $\ell$ contained in the unbounded face of $G$, and the two new edges $\{w,v\}$ and $\{w,v^{\prime}\}$ (both weighted by 1), where $v^{\prime}$ is the mirror image of $v$ (see the picture on the right in Figure 4). Then $\overline{G}$ is a plane, bipartite, weighted symmetric graph, to which part (a) of the theorem can be applied.
@@
If $v$ is white, it follows that $w$ is black, and (since $w$ is also a $b_{i}$-type vertex) the cutting operation at $w$ prescribed by the factorization theorem occurs above $w$. Therefore, $\overline{G}^{+}$ is precisely $G^{+}$. Furthermore, in $\overline{G}^{-}$ vertex $w$ has degree one, and after removing the edge $\{u,v^{\prime}\}$ that is forced to match it in every perfect matching, the resulting graph (which has the same matching generating function as $\overline{G}^{-}$, as the removed edge has weight 1) is precisely $G^{-}\setminus v^{\prime}$. Thus part (a) of the theorem applied to $\overline{G}$ yields
@@
Draw the triangular lattice so that one family of lattice lines is horizontal, and let $H$ be a non-degenerate (i.e., with all sides of positive length) hexagonal region on this lattice, symmetric about a vertical symmetry axis666 We will be interested in symmetric (or nearly-symmetric) regions obtained from $H$ by making in it some holes, and since the symmetry of such regions seems to be more readily appreciated visually when the symmetry axis is vertical, we adopt this point of view for the second result in this section. $\ell$. Let $R_{0}$ be the region obtained from $H$ by removing an arbitrary collection $\mathcal{C}$} of unit triangles, so that $\mathcal{C}$} is symmetric about $\ell$ and contains no unit triangles touching the southern or southeastern edge of $H$. Define $R$ to be the region obtained from $R_{0}$ by translating all unit triangles in $\mathcal{C}$} one unit in the southeastern lattice direction. We call the region $R$ a nearly symmetric hexagon with holes (an example is shown in the picture on the left in Figure 5). Therefore, while the outer, hexagonal boundary of the region $R$ is symmetric with respect to the original symmetry axis $\ell$, the holes in $R$ are symmetric about the line $\ell^{\prime}$, the translation of $\ell$ half a unit to the right.
Let $G$ be the planar dual graph of $R$. Even though $G$ is not symmetric about $\ell^{\prime}$, let us consider the graphs $G^{+}$ and $G^{-}$ obtained from it by rotating it 90 degrees counterclockwise and performing cutting operations around the vertices on $\ell^{\prime}$ as prescribed by the factorization theorem (assuming, in particular, that after the rotation the leftmost vertex on $\ell^{\prime}$ is white). In addition to these two graphs, consider also the graphs $\widehat{G}^{+}$ and $\widehat{G}^{-}$ obtained from $G$ by performing cutting operations around the vertices on $\ell^{\prime}$ as prescribed by the factorization theorem, but assuming this time that after the rotation the leftmost vertex on $\ell^{\prime}$ is black. Define $^{+}, RR^{-}$, $\widehat{R}^{+}$ and $\widehat{R}^{-}$ to be the lattice regions whose planar duals are the graphs $^{+}, GG^{-}$, $\widehat{G}^{+}$ and $\widehat{G}^{-}$, respectively777 If $R$ is the region on the triangular lattice whose planar dual is the graph $G$, and $G$ has weights on its edges, the lozenge position in $R$ corresponding to any given edge of $G$ comes weighted with the weight of that edge..
@@
Let $R$ be a nearly symmetric hexagon with holes, and assume that the symmetric region $R_{0}$ from which $R$ was constructed888 I.e. the region obtained from $R$ by translating all the holes one unit lattice segment in the northwest direction. admits at least one lozenge tiling. Then the number of lozenge tilings999 We denote the matching generating function of the lozenge tilings of a region $R$ (in which lozenge positions may carry weights) by $\operatorname{M}(R)$, because these lozenge tilings can be identified with perfect matchings of the (weighted) planar dual graph of the region $R$. of $R$ is given by
@@
[…]
@@
[…]